# Large Language Models Can Learn Temporal Reasoning

### v2

@@ Appendix A TGQA Dataset Statistics
Table 6: We use a global mapping for entity names from GPT-3.5 to avoid corpus leakage.
@@
Table 7: We integrate the necessary mathematics in context as external knowledge for temporal reasoning.
Appendix B Experiment Details
In this section, we present more experiment details for further research. We include data augmentation examples, model versions and evaluation tasks & metrics.
@@ Appendix B Experiment Details
CoT (symbolic): The duration for each event can be
@@
CoT (symbolic): The duration for each event can be
@@
Table 8: We propose special data augmentation strategies for symbolic reasoning over (temporal) graphs.
Data Augmentation Examples  We show with an example our data augmentation strategies in Table 8. For TGs with relation synonym replacement or irrelevant edges removal, there is no need for change on Graph-based QAs. For TGs with entity names/times mapping, we also need to change with the corresponding entity names/times in Graph-based QAs to ensure the consistency.
Model Versions  All the model versions of LLMs used in our experiments are listed below:
  • Llama2-7B (meta-llama/Llama-2-7b-hf)
  • Llama2-13B (meta-llama/Llama-2-13b-hf)
  • Llama2-70B (meta-llama/Llama-2-70b-hf)
  • GPT-3.5 (gpt-3.5-turbo)
  • GPT-4 (gpt-4-1106-preview)
For the Llama2 family, all the model weights can be downloaded from the platform of Hugging Face. For the GPT models, all the model weights can be accessed through the OpenAI APIs.
Evaluation Tasks & Metrics  In addition to TGQA, we consider the existing datasets TimeQA  (Chen et al., 2021) and TempReason  (Tan et al., 2023a), which are constructed using Wikipedia articles, excerpts, or summaries. To be specific, TimeQA  (Chen et al., 2021) contains two difficulty levels (see Table 9). The easy-level split tends to be information extraction while the hard-level split involves understanding the relation between different temporal expressions. TempReason  (Tan et al., 2023a) contains three levels (L1: Time-Time Relation, L2: Time-Event Relation, L3: Event-Event Relation) and three settings (OBQA: open-book QA, CBQA: close-book QA, ReasonQA: facts-based QA) (see Table 10). We observe that some stories in TempReason are incomplete which partially leads to the low accuracy of LLMs. We evaluate our framework on all the datasets with the metrics of token-level F1, exact match (EM) and perplexity-based accuracy (Acc). Besides choosing F1 and EM, which are two basic metrics for span-based QA tasks, we consider Acc for LLM evaluations.
@@
Compared with other LMs, LLMs are pre-trained on a wide variety of tasks, making them more generalizable at the cost of less conforming to the required format. This property actually hurts their performance under the commonly used metrics such as EM and F1. To solve this problem, we select from a candidate set the answer with the lowest perplexity from LLMs as their prediction. For questions with multiple correct answers, we follow the strategy proposed in  (Chen et al., 2021) to get the best F1, EM and Acc among them. Since there are multiple categories of questions in TGQA, we first calculate the metrics for each category, and then use the average as final scores to mitigate question category imbalance.
@@
${}^{{\ddagger}*}$ (0-shot SP) 0.801 0.689 0.920 0.620 0.930 0.940 0.758 0.575 0.737 0.776
$*{}^{*}$ (1-shot SP) 0.820 0.660 0.870 0.630 0.910 0.990 0.760 0.610 0.700 0.772
${}^{{\ddagger}*}$ (0-shot CoT) 0.950 0.692 0.901 0.881 0.919 0.980 0.778 0.647 0.858 0.848
$*{}^{*}$ (1-shot CoT) 0.930 0.758 0.903 0.869 0.892 0.920 0.758 0.667 0.680 0.821
@@
Llama2-13B (1-shot SFT - TG + SR(bs)) 0.947 0.732 0.747 0.825 0.931 0.987 0.495 0.466 0.649 0.753
Llama2-13B (1-shot SFT - TG + SR(bs + aug)) 0.931 0.710 0.729 0.860 0.927 0.977 0.627 0.534 0.739 0.782
Llama2-13B (1-shot SFT - TG + EK + SR(bs + aug)) 0.944 0.775 0.729 0.849 0.924 0.980 0.783 0.506 0.681 0.797
@@
We explain the categories of errors in human evaluation on CoT generation here (see Table 3). T1 means that LLM uses wrong information such as wrong start/end time during reasoning. T2 means that LLM makes logical errors, e.g., after determining "Event A ends before Event B starts", considering the statement as true that "Event A was still happening when Event B starts". T3 means that LLM makes errors on external knowledge, e.g., claiming that "the date 1978 is after 1983". T4 means that there exists error in the extracted TG which leads to the wrong conclusion.
@@ Appendix C Example Prompts
In this section, we show example prompts used in our framework. To be specific, Table 12 shows an example of the graph-based story generation in TGQA; Table 13 presents an example of the automatic story-temporal graph alignment verification in TGQA; Table 14 provides an example of the temporal info identification in TimeQA; Table 15 illustrates an example of the graph construction in TimeQA; Table 16 gives an example of the automatic temporal graph-QA alignment verification in TimeQA; Table 17 depicts an illustrative case of the CoTs bootstrapping in TGQA.
Appendix D Fine-grained Results
We provide fine-grained results on TGQA in Table 11. In TGQA, there are nine types of questions as explained in Appendix A. We show all the models with different strategies here. Zero-shot performance of LLMs is also considered to investigate the effect of in-context examples. Note that we do not include zero-shot performance of the Llama2 family since it is not fine-tuned on instructions, i.e., we can not obtain valid zero-shot learning results. For zero-shot learning results, the format of generation is not guaranteed. The original accuracy is very low since rule-based parser cannot handle it. Instead, we use GPT-3.5 as parser to extract the answer from the generation. Also, we only adopt one example in context due to context length limitation.
It can be seen that in-context examples are not necessary for powerful pre-trained models such as GPT-3.5 and GPT-4. Undoubtedly, GPT-4 obtains the best in-context learning performance on most categories. However, for some categories, the best performance of GPT-4 is around 0.6 or 0.7, which is not sufficiently well. On the other hand, we show that our strategies improve the performance step by step. Note that our model achieves the best overall performance by improving all the categories. To be specific, symbolic reasoning brings better performance for most categories. Furthermore, data augmentation and external knowledge further enhance the capability of the model while preserving the existing good performance on those categories.
@@ Appendix D Fine-grained Results
Table 13: Example of the automatic story-temporal graph alignment verification in TGQA.
@@
Table 15: Example of the graph construction in TimeQA.
@@
Table 16: Example of the automatic temporal graph-QA alignment verification in TimeQA.
@@
Table 17: Example of the CoTs bootstrapping in TGQA.

### v3

@@ Appendix B Experiment Details
Table 6: We use a global mapping for entity names from GPT-3.5 to avoid data leakage.
@@
Table 7: We integrate the necessary mathematics and commonsense in context as external knowledge for TR.
@@
CoT: The duration for each event can be
@@
CoT: The duration for each event can be
@@
Table 8: We propose several graph data augmentation strategies for reasoning over temporal knowledge graph. The original TG is shown in Table 1. We highlight the changed information.
Graph Data Augmentation.  We show with an example our graph data augmentation strategies in Table 8. For TGs with relation synonym replacement or irrelevant edges removal, there is no need for change on graph-based QAs. To contrust, for TGs with entity names/times mapping, we need to change with the corresponding entity names/times in graph-based QAs to ensure the consistency.
Evaluation Tasks & Metrics.  Besides TGQA, we consider in experiments the two existing datasets TimeQA  (Chen et al., 2021) and TempReason  (Tan et al., 2023a). Specifically, TimeQA contains two difficulty levels (Table 9). The easy-level split tends to be the information extraction task while the hard-level split involves understanding the relation between different temporal expressions. On the other hand, TempReason contains three levels (L1: Time-Time Relation, L2: Time-Event Relation, L3: Event-Event Relation) and three settings (OBQA: open-book QA, CBQA: close-book QA, ReasonQA: facts-based QA) (Table 10). We observed that some stories in TempReason are incomplete which partially leads to the low accuracy of LLMs.
@@
We evaluate our framework on all the datasets with the metrics of token-level F1, exact match (EM) and perplexity-based accuracy (Acc). F1 and EM are two basic metrics for span-based QA tasks. However, the free-form prediction of LLMs might hurt their performance under these generation-based metrics. To solve this problem, we introduce perplexity-based accuracy, i.e., selecting from a candidate set the final answer with the lowest perplexity. For questions with multiple correct answers, we follow the strategy proposed in  (Chen et al., 2021) to get the best result among them. Since there are multiple question categories in TGQA, we first calculate the metrics for each category, and then use the average as final scores to mitigate question category imbalance.
@@
GPT-4‡∗ (0-shot SP) 0.801 0.689 0.920 0.620 0.930 0.940 0.758 0.575 0.737 0.776
GPT-4 (1-shot SP) 0.820 0.660 0.870 0.630 0.910 0.990 0.760 0.610 0.700 0.772
GPT-4‡∗ (0-shot CoT) 0.950 0.692 0.901 0.881 0.919 0.980 0.778 0.647 0.858 0.848
GPT-4 (1-shot CoT) 0.930 0.758 0.903 0.869 0.892 0.920 0.758 0.667 0.680 0.821
@@
Llama2-13B (1-shot SFT - TG + CoT(bs)) 0.947 0.732 0.747 0.825 0.931 0.987 0.495 0.466 0.649 0.753
Llama2-13B (1-shot SFT - TG + CoT(bs + aug)) 0.931 0.710 0.729 0.860 0.927 0.977 0.627 0.534 0.739 0.782
Llama2-13B (1-shot SFT - TG + EK + CoT(bs + aug)) 0.944 0.775 0.729 0.849 0.924 0.980 0.783 0.506 0.681 0.797
@@
Human evaluation on CoT generation found out four types of error in total (Table 3). T1 means that LLM uses wrong information such as wrong start/end time during reasoning. T2 suggests that LLM makes logical errors, e.g., mentioning "Event A ends before Event B starts" in CoT but determining the statement "Event A was still happening when Event B starts" to be true. T3 denotes that LLM makes errors on external knowledge, e.g., claiming that "the date 1978 is after 1983". T4 indicates that there exists errors in the extracted TG that lead to the wrong conclusion.
Candidate Answer Generation. We involve candidate answers in the calculation of Acc. For TempReason, the authors provide negative answers since they perform time-sensitive reinforcement learning. That is, they use the score of the correct answers and wrong answers from the language model as reward to further finetune the model parameter. For TimeQA, we generate the candidates in the following way. Given a story, there are multiple related questions which share the same subject/object entity and relation. The correct answer changes with the query time. For example, "What position did George Washington hold in 1777/1790/1799?" Answer: "Commander in Chief/Presidency/Chancellor". We collect the answers of these related questions as candidates. More generally, if there are no such related questions, we will use all the entities in the corresponding TG as candidates.
Model Versions.  The versions of the LLMs used in our experiments are listed below. For the Llama2 family, all the model weights can be downloaded from the platform of Hugging Face. For the GPT models, all the model weights can be accessed through the OpenAI APIs.
  • Llama2-7B (meta-llama/Llama-2-7b-hf)
  • Llama2-13B (meta-llama/Llama-2-13b-hf)
  • Llama2-70B (meta-llama/Llama-2-70b-hf)
  • GPT-3.5 (gpt-3.5-turbo)
  • GPT-4 (gpt-4-1106-preview)
@@ Appendix C Example Prompts
In this section, we show example prompts used in our framework. Specifically, Table 12 shows an example of the graph-based story generation in TGQA; Table 13 presents an example of the automatic story-temporal graph alignment verification in TGQA; Table 14 provides an example of the temporal info identification in TimeQA; Table 15 illustrates an example of the graph construction in TimeQA; Table 16 gives an example of the automatic temporal graph-QA alignment verification in TimeQA; Table 17 depicts an illustrative case of the CoTs bootstrapping in TGQA.
Appendix D Fine-grained Results of TGQA
We provide the fine-grained results of TGQA in Table 11. In TGQA, there are nine types of questions as explained in Appendix A. We show all the models with different strategies. Zero-shot performance is also considered to investigate the effect of in-context examples. Note that we do not include zero-shot performance of the Llama2 family since it is not fine-tuned on instructions, i.e., we can not obtain valid zero-shot learning results. For zero-shot learning results, the format of generation is not guaranteed. The original accuracy is very low since rule-based parser cannot handle it. Instead, we use GPT-3.5 as parser to extract the answer from the generation. Also, we only adopt one example in context due to context length limitation.
It can be seen that in-context examples are not necessary for powerful pre-trained models such as GPT-3.5 and GPT-4. Undoubtedly, GPT-4 obtains the best in-context learning performance on most categories. However, for some categories, the best performance of GPT-4 is around 0.6 or 0.7, which is not sufficiently well. On the other hand, we show that our strategies improve the performance step by step. Note that our model achieves the best overall performance by improving all the categories. Specifically, deliberate reasoning with CoT bootstrapping brings better performance for most categories. Furthermore, graph data augmentation and external knowledge further enhance the capability of the model while preserving the existing good performance on those categories.
@@ Appendix D Fine-grained Results of TGQA
Table 13: Example of the story-temporal graph alignment verification in TGQA.
@@
Table 15: Example of the temporal graph construction in TimeQA.
@@
Table 16: Example of the temporal graph-QA alignment verification in TimeQA.
@@
Table 17: Example of the CoT bootstrapping in TGQA.